OpenStudy (anonymous):

Help!?!?!?!? 13. What is a quartic function with only the two real zeroes given? x = –1 and x = –3 (1 point) y = x4 –4x3 – 4x2 – 4x – 3 y = –x4 + 4x3 + 4x2 + 4x + 3 y = x4 +4x3 + 3x2 + 4x – 4 y = x4 + 4x3 +4x2 + 4x + 3

OpenStudy (anonymous):

13. What is a quartic function with only the two real zeroes given? x = –1 and x = –3 (1 point)

OpenStudy (anonymous):

can sombody help me pleaaz

OpenStudy (raden):

one of way to solve it, by substitute back the roots to original equation u can do one by one from options and u will found the best anwer

OpenStudy (raden):

which the options, when u subs x=-1 and x=-3 to get y=0

OpenStudy (anonymous):

so it A

OpenStudy (raden):

let's see is A correct or no :p

OpenStudy (raden):

y = x^4 –4x^3 – 4x^2 – 4x – 3, first subs x=-1 y= (-1)^4 -4(-1)^3 - 4(-1)^2 -4(-1) - 3 y= 1 + 4 - 4 + 4 - 3 y=2 because y not equal 0, so A incorrect try again :p

OpenStudy (anonymous):

its not b

OpenStudy (raden):

can u show ur job, how do u know the answer is B or other

OpenStudy (anonymous):

its not b because i t = -6

OpenStudy (anonymous):

is it d

OpenStudy (raden):

yea, go to c or d

OpenStudy (raden):

how comes, show to me ur calc

OpenStudy (anonymous):

is it d

OpenStudy (anonymous):

i need the answer fast pleazz

OpenStudy (raden):

hahaha... ksian cucu atok well, the answer is D.. u r correct

OpenStudy (raden):

any else

OpenStudy (anonymous):

10. Find all the zeroes of the equation. x3 + x2 = –3x – 3 (1 point)

OpenStudy (raden):

first move side all right side to left side, it can be x^3 + x^2 + 3x + 3 = 0 do trials, for x=-1 would satisfies as the root it means (x+1) is one of the factor. from here u can found the other factors by use synthetic division

OpenStudy (anonymous):

i substitued -1 and got 0=0

OpenStudy (anonymous):

the answer has to be a or c

OpenStudy (raden):